OpenStudy (anonymous):

Solve: 4(x − 1) = 2(x + 2) + 6 14 8 7 28

jimthompson5910 (jim_thompson5910):

Start off by distributing

OpenStudy (anonymous):

i did, 4x-4=2x+4+6

jimthompson5910 (jim_thompson5910):

now combine like terms

OpenStudy (anonymous):

6x - 4 = 10?

jimthompson5910 (jim_thompson5910):

no, only combine like terms on the same side

OpenStudy (anonymous):

thats confusing

jimthompson5910 (jim_thompson5910):

left side: are there any like terms in 4x-4 ?

OpenStudy (anonymous):

O.o

jimthompson5910 (jim_thompson5910):

can we combine 4x and -4 ?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

so there are no like terms on the left side

OpenStudy (anonymous):

no there isnt

jimthompson5910 (jim_thompson5910):

what about the right side?

OpenStudy (anonymous):

4 and 6

jimthompson5910 (jim_thompson5910):

so they add to 10

jimthompson5910 (jim_thompson5910):

which means 4x-4=2x+4+6 becomes 4x-4=2x+10

OpenStudy (anonymous):

yes ik

jimthompson5910 (jim_thompson5910):

ok tell me what you get when you solve

OpenStudy (anonymous):

idk what to do now

jimthompson5910 (jim_thompson5910):

now get all the x terms to one side and get everything else to the other side

OpenStudy (anonymous):

4x-2x=-4+10?

jimthompson5910 (jim_thompson5910):

close, but it should be 4x - 2x = 10+4

OpenStudy (anonymous):

2x=14? x=7?

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

yw